Consider the reaction for the decomposition of hydrogen disulfide:

2H2S(g)⇌2H2(g)+S2(g), Kc = 1.67×10−7 at 800∘C

A 0.500 L reaction vessel initially contains 0.175 mol of H2S and 6.25×10−2 mol of H2 at 800∘C.

Find the equilibrium concentration of  [S2].
Answer #1

Volume of the vessel = 0.500 L

Initial concentration of H2S (g) = 0.175/0.500 = 0.35 M

Initial concentration of H2 (g) = 6.25×10−2 /0.500 = 0.125 M

                                                                     2H2S(g)    2H2(g) + S2(g)

Initial concentration                                       0.35                0.125         0

Change in concentration                                - 2X                  2X           X

Equilibrium concentration                          (0.35 - 2X)       (0.125+2X)   X

Equilibrium constant,

Kc = [H2]2[S2]/[H2S]2

or, 1.67 x 10-7 = (0.125+2X)2(X)/(0.35 - 2X)2

or, 1.67 x 10-7 = (0.125)2(X)/(0.35 )2     [(0.35 - 2X) 0.35 as 2X << 0.35 and (0.125+2X) 0.125 as 2X << 0.125]

or, 0.015625X = 0.0000000204575

or, X = 1.31 x 10-6 M

Therefore, the equilibrium concentration of S2(g), [S2] = 1.31 x 10-6 M